OpenStudy (anonymous):

At Harold's Haberdashery, men's overcoats cost x dollars and women's overcoats cost $2 more. If a costume designer buys 10 men's coats and 14 women's coats, which of the following represents the total value of the purchase?

OpenStudy (anonymous):

10x+14(x+2)=y 10x+14x+28=y 24x+28=y.

OpenStudy (anonymous):

X being the cost of the men's overcoats, and Y being the total cost of the purchase.

OpenStudy (anonymous):

thank you